express the equation in exponential form. please show the step of hot to get the correct answer without skipping any steps.
a) In(x+9)=6
b) In(x-9)=8

Answer by lwsshak3(11628)

express the equation in exponential form. please show the step of hot to get the correct answer without skipping any steps.
exponential form: base(e) raised to log of number(6)=number(x+9)
a) In(x+9)=6
exponential form: base(e) raised to log of number(6)=number(x+9)
e^6=x+9
x=e^6-9
..
b) In(x-9)=8
base(e) raised to log of number(8)=number(x-9)
e^8=x-9
x=e^8+9